OpenStudy (anonymous):

x,y and z are non-negative integers where \(\large xy+3x+2y=2\) and \(\large yz+4y+3z=52\) Find the value of \(\large x+y+z+xy+yz+xz+xyz\)

OpenStudy (anonymous):

Any idea?

OpenStudy (anonymous):

I would start out substituting

OpenStudy (anonymous):

yeah, I think too

OpenStudy (anonymous):

So what does xy equal?

OpenStudy (anonymous):

\(\large xy=2-3x-2y\)

OpenStudy (anonymous):

Exactly same for yz then plug into equation I am working on it but am stuck at xyz

OpenStudy (anonymous):

Unless you convert and multiply. Say x = (-xy+2y-2)/3 then do the same for y and z.

OpenStudy (anonymous):

xy+3x+2y=2 x(y+3)=2(1-y) x=(1-y)/(y+3) Since x,y,z are non negative integers only solution of x and y is when y=1 and x=0

OpenStudy (anonymous):

and when y=1, 1*z+4*1+3z=52 4z=48 z=12

OpenStudy (anonymous):

so, x+y+z+xy+yz+xz+xyz=25

OpenStudy (anonymous):

Thanks @sauravshakya . Your solution is concise and nice And Here is my solution We have \(\large \begin{cases}\large xy+3x+2y=2\ \ \ (1)\\ \large yz+4y+3z=52\ \ \ (2) \end{cases} \Rightarrow \begin{cases} y=\frac{2-3x}{x+2}\\ y=\frac{52-3z}{z+4} \end{cases}\) \(\large \Rightarrow \frac{2-3x}{x+2}=\frac{52-3z}{z+4} \Leftrightarrow z-8x-12=0 \Rightarrow z=8x+12\ \ \ (3)\) Now, substituting \(\large (3)\) into \(\large(2)\) give \(\large y(8x+12)+4y+3(8x+12)=52\\\large \Leftrightarrow 2xy+3x+2y=2\ \ \ (4)\) Then, subtract \(\large (1)\) from \(\large (4)\) got:\(\large xy=0\) \(\large \Rightarrow x=0\ Or\ y=0\) \(\Large \color{red}{*x=0} \Rightarrow y=1\ \ and \ z=12\) (satisfy x,y and z are non-negative integers) \(\Large \color{red}{*y=0} \Rightarrow x= \frac{2}{3}\) (not satify ) Thus, \(\large x+y+z+xy+yz+xz+xyz = (x+1)(y+1)(z+1)-1=25\)